Find the real zeros of each polynomial function by factoring. The number in parentheses to the right of each polynomial indicates the number of real zeros of the given polynomial function. (Enter your answers as a comma-separated list.)
P(x) = x5 − 17x3 + 16x (5)
x =

Answer by mananth(16946)


x5-17x3+16x=0

⇒x5-17x3+16x=0

⇒x(x4-17x2+16)=0

⇒x(x4-x2-16x2+16)=0

⇒x(x2(x2-1)-16(x2-1))=0

⇒x(x2-1)(x2-16)=0

x(x+1)(x-1)(x2-4)=0

x(x+1)(x-1)(x+2)(x-2)=0

x=0or(x-1)=0or(x+1)=0or(x-4)=0or(x+4)=0